OpenStudy (anonymous):

Solve. 2(n - 3) + 5 = 3(n - 1) Solve. 2(n - 3) + 5 = 3(n - 1) @Mathematics

OpenStudy (alfie):

2n-6 + 5 = 3n -3. 2n - 3n -1 = -3 -n = -2 n = 2